Lithium hydroxide can be used to remove carbon dioxide exhaled by astronauts. The reaction  involves lithium hydroxide reacting with carbon dioxide to produce lithium hydrogen carbonate.  Each cartridge that is used has a lithium hydroxide mass of 3.00 x 104 g. If the process for  removing carbon dioxide from the air has a percentage yield of 83.0 %, what mass of carbon  dioxide is removed by each canister of lithium hydroxide.




1
Expert's answer
2021-03-22T06:39:17-0400

2LiOH + CO2 "\to" Li2CO3 + H2O

Moles of Lithium Hydroxide = 3×104/23.95 = 12.52×102

2 moles of LiOH react with 1 mole of CO2

So ,

Moles of Carbon dioxide = 6.26×102

Mass of carbon dioxide = 6.26×102 × 44.01 = 2.75×104 g

Percentage yield = (Actual yield/theoretical yield)×100

83 =( Actual yield / 2.75×104 g)×100

Actual yield =( 2.75×104 ×83) /100

= 2.28×104

Removed = theoretical yield - actual yield

= 2.75×104 --2.28×104

= 4.7×103 g